Find the slope of the line:
y=6x+4

Found 2 solutions by Positive_EV, General_Lee87:
Answer by Positive_EV(69)

Whenever any line is in the form y = mx + b for any constants m and b, the line is said to be in "slope-intercept" form. For any line in slope-intercept form, m is equal to the slope of the line, and b is equal to the y-intercept of the line.

for y = 6x+4, m = 6 and b = 4, so the slope is 6.

Answer by General_Lee87(58)

Origional Question: Find the slope of the line: y=6x+4
Answer/Worked Out Problem:
y=6x+4
Slope is before x in a y-intercept form equation
Therefore: m=6 (m meaning slope)
I hope this helped.
God bless you and have a wonderful and prosperous New Year!